Becky Lewis estimates that she will need $35,000 in five years for a down payment on a new home. What amount would she have to invest today at 2.1% interest, compounded monthly, in order to reach her goal?

  • $31,514.25

  • $31,671.95

  • $27,012.21

  • $34,695.35
Answer #1

We use the formula:
A=P(1+r/12)^12n
where
A=future value
P=present value
r=rate of interest
n=time period.

35000=P*(1+0.021/12)^(12*5)

P=35000/(1+0.021/12)^(12*5)

=35000*0.900407147

=$31514.25(Approx).